OpenStudy (anonymous):

Let g be the function given by g(x)=(x+4)/√7x-5. find and simplify 4g(3)+2[g(2)]^2. Can someone please help me with this question. would like to see the steps on how to solve please thanks again.

OpenStudy (anonymous):

Ok so this is how our equation will look like \[4 ( \frac{(3)+4}{\sqrt{7(3)-5}})+2(\frac{(2)+4}{\sqrt{7(2)-5}})^2\]

OpenStudy (campbell_st):

start by finding the value of g(3) then also find the value of g(2) then when you have the values multiply g(3) by 4 and add g(2) square multiplied by 2 hope it makes sense

OpenStudy (anonymous):

Thank you BlackLabel for your help.

OpenStudy (anonymous):

Thanks Campbell _st :)